OpenStudy (buttermequeasy):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x

OpenStudy (anonymous):

if it's parallel to y = -3x -4, it means it has the SAME slope. -3 = m, where m is the slope. (of the equation y = mx +b) so all ya gotta do is plug all the values (-3 = m, and your point (2,6)) together to find a new line parallel to that one: y - y1 = m(x - x1) y - 6 = -3(x - 2) y - 6 = -3x + 6 y = -3x